Josh believes the Spanish club students at his school have an unfair advantage in being assigned to the Spanish class they request. He asked 500 students at his school the following questions: "Are you in the Spanish club?" and "Did you get the Spanish class you requested?" The results are shown in the table below.

Spanish Club Not in Spanish Club Total
Received Spanish class requested 265 100 365
Did not get Spanish class requested 70 65 135
Total 335 165 500
Help Josh determine if all students at his school have an equal opportunity to get the Spanish class they requested. Show your work and explain your process for determining the fairness of the class assignment process.

Answer by richwmiller(17219)

sc nsc total
yes 265 100 365
no 70 65 135
tot 335 165 500
65/165=.3939 nsc 39.39% NSC didn't receive their choice
70/335=.20 sc 20% of the spanish club didn't receive their choice
135/500=.27 school 27% of the school didn't receive their choice
100/165 nsc .6061 61% NSC DID receive their choice
265/335 sc .79 79 % of the Spanish club DID receive their choice
365/500 school 73% school DID receive their choice